If X is the random variable model ‘the number on a biased die’, and the probability distribution function of X is as shown below,


1
Expert's answer
2021-10-22T16:44:28-0400

As here the question is incomplete, so suggested the solution based on the given terms.

Given,

Random variable on biased die = X

p(x) = P(X=x) for all the value of x

Now, the variance of a random variable X

"\Rightarrow V(X)= E[(X-\mu)^2]"

For the case of the die toss,

Expected values "E(X)=\Sigma_{x\in X}(xP(x))"

Now, for the calculation of variance,

"E(X^2)=\Sigma x^2 P(x)"

"Var(X)=E(X^2)-E^2(X)"